OpenStudy (anonymous):

solve by the substution method y=x^2-5x-6 y=x^2+2x+8

OpenStudy (mertsj):

\[x^2-5x-6=x^2+2x+8\] \[-7x=14\] \[x=-2\]

OpenStudy (anonymous):

its has to be in a form like y=x^2+9x+9 y=x^2+2x-19 The answer to this one is (-4,-11)

OpenStudy (anonymous):